Use a 2 step Eulers method to approximate y(1.2), of the solution of the initial-value problem y = 1 – 2x2 – 2y, y(1) = 4.
Answer #1

The Eulers method states that yn+1 = yn +hf (In, Yn), where In+1 = In + h. We have that h = 1 , 10 = 1, yo = 4, f(x, y) = –2The Eulers method states that Yn+1 = yn +h: f(In, yn), where Xn+1 = xn + We have that h = 1 , X0 = 1, yo = 4, f(x, y) = –2x2